Evaluate the expotential function below for x=-4.
J(x)=(1/8)^-4
a. 1/4096
b.8
c.1/512
d.4096
e.512

Found 2 solutions by jorel1380, ccs2011:
Answer by jorel1380(3719)

(1/8)^-4=1/(1/8)4=1/(1/4096)=4096..

Answer by ccs2011(207)